# Searching Weighted Barbell Graphs with Laplacian and Adjacency Quantum Walks

### v1

@@ I Introduction
In discrete space, which can be modeled as a graph of $N$ vertices and edges, the particle can only be at the vertices, and it can only jump along edges. This is known as a continuous-time quantum walk, and is the quantum analogue of a continuous-time random walk [2]. The vertices correspond to computational basis states ${\left|1\right\rangle},{\left|2\right\rangle},\dots,{\left|N\right\rangle}$, and the state ${\left|\psi\right\rangle}$ of the particle is generally a superposition over the vertices. In the Hamiltonian, the continuous-space Laplace operator $\nabla^{2}$ is replaced by the discrete Laplacian
@@ II Search Algorithm
As in [16], we can reduce the dimension of the space using the symmetry of the barbell graph. As labeled in Fig. 3, the barbell graph has five types of vertices, $a$, $b$, $c$, $d$, and $e$. Each vertex of the same type evolves identically, e.g., as the system evolves, the amplitude at each of the $b$ vertices is the same. So, the system evolves in a 5D subspace spanned by the orthonormal basis consisting of uniform superpositions over each type of vertex:
@@ III.1 Small Weights
Then, linear combinations of ${\left|a\right\rangle}$, ${\left|b\right\rangle}$, and ${\left|e\right\rangle}$ are eigenvectors of $H_{L}{(0)}+H_{L}^{(1)}$, i.e.,
@@ V Two-Stage Algorithm
The success probability $p_{a}blue{(\Delta t)}$ reaches its maximum value of
@@
The peak success probability of $0.996$ is independent of $N$, as verified by Fig. 10, where we have plotted the two-stage search algorithm for various values of $N$. They all reach the same height of $0.966$. This figure also confirms the runtime expression (30), since the solid black curve peaks at time $4.610\sqrt{1024}\approx 147.5$ (as also seen in Fig. 7), the dashed red curve peaks at time $4.610\sqrt{2048}\approx 208.6$, and the dotted green curve peaks at time $4.610\sqrt{4096}\approx 295.0$.
@@ VI Conclusion
Previous research exploring search on the weighted truncated simplex lattice and linked complete graphs have shown that the weights between the cliques could be tuned to improve the runtime and success probability of quantum walk search algorithms by allowing probability to be freed from cliques. In these, the Laplacian and adjacency quantum walks searched in the same manner. In this paper, we explored the most restrictive means to join two cliques by investigating the weighted barbell graph of $N$ vertices, which has a single bridge of weight $w$. The Laplacian and adjacency quantum walks generally search this graph differently. For the Laplacian quantum walk, the weighted edge makes no difference; the evolution is the same regardless of the weight, where the probability in confined in each clique, and so the success probability reaches a maximum value of $1/2$ at time $\pi\sqrt{N}/2\sqrt{2}$. That is, the single bridge is too restrictive to affect the Laplacian quantum walk. For the adjacency quantum walk, most weights also make no difference, with the system evolving just like the Laplacian quantum walk. When $w=N/2$, however, the adjacency quantum walk does free some of the probability that was trapped in the unmarked clique, allowing the success probability to reach a value of 0.820 at time $2.518\sqrt{N}$, where the coefficient is the solution to a transcendental equation. Evolving beyond this, to a time of $3.265\sqrt{N}$, decreases the success probability, but causes 99.6% of the total probability to be in the marked clique. Then, the weight can be changed, for example to $w=1$, and the system will evolve to focus the success probability to the marked vertex, which takes an additional time of $1.345\sqrt{N}$. Altogether, this two-stage algorithm takes $4.610\sqrt{N}$ time to reach a success probability of $0.996$. Although this two-stage algorithm takes longer than search on the unweighted graph with repetitions, it could be useful in a scenario where it is expensive to repeat an algorithm, beyond the cost of querying the oracle.

### v2

@@ I Introduction
In discrete space, which can be modeled as an undirected graph of $N$ vertices and edges, the particle can only be at the vertices, and it can only jump along edges. This is known as a continuous-time quantum walk, and is the quantum analogue of a continuous-time random walk [2]. The vertices correspond to computational basis states ${\left|1\right\rangle},{\left|2\right\rangle},\dots,{\left|N\right\rangle}$, and the state ${\left|\psi\right\rangle}$ of the particle is generally a superposition over the vertices. In the Hamiltonian, the continuous-space Laplace operator $\nabla^{2}$ is replaced by the discrete Laplacian
@@ II Search Algorithm
As in [16], we can reduce the dimension of the space using the symmetry of the barbell graph. As labeled in Fig. 3, the barbell graph has five types of vertices, $a$, $b$, $c$, $d$, and $e$, where $a$ is the marked vertex, $c$ and $d$ are the endpoints of the bridge ($c$ in the marked clique and $d$ in the unmarked one), $b$ are the vertices in the marked clique except for $a$ and $c$, and $e$ are the vertices in the unmarked clique, except for $d$. Each vertex of the same type evolves identically, e.g., as the system evolves, the amplitude at each of the $b$ vertices is the same. So, the system evolves in a 5D subspace spanned by the orthonormal basis consisting of uniform superpositions over each type of vertex:
@@ III.1 Small Weights
Then, linear combinations of ${\left|a\right\rangle}$, ${\left|b\right\rangle}$, and ${\left|e\right\rangle}$ are eigenvectors of $H_{L}^{(0)}+H_{L}^{(1)}$, i.e.,
@@
On the other hand, when $\gamma$ is away from its critical value of $\gamma_{c}=2/N$ (13), then an examination of the eigenvectors and eigenvalues of $H_{L}^{(0)}$ in (12) reveals that ${\left|a\right\rangle}$ is no longer degenerate with ${\left|b\right\rangle}$ and ${\left|d\right\rangle}$. As a result, the eigenvectors and eigenvalues of $H_{L}^{(0)}+H_{L}^{(1)}$ are no longer those from (15). Instead, ${\left|b\right\rangle}$ and ${\left|e\right\rangle}$ are each eigenvectors of $H_{L}^{(0)}+H_{L}^{(1)}$ with eigenvalue 0, and since the initial state ${\left|\psi(0)\right\rangle}\approx({\left|b\right\rangle}+{\left|e\right\rangle})/\sqrt{2}$, the system asymptotically stays in this initial state, and the quantum walk does not build up probability at the marked vertex.
@@ V Two-Stage Algorithm
The success probability $p_{a}{(\Delta t)}$ reaches its maximum value of
@@
The peak success probability of $0.996$ is independent of $N$, as verified by Fig. 10, where we have plotted the two-stage search algorithm for various values of $N$. They all reach the same height of $0.996$. This figure also confirms the runtime expression (30), since the solid black curve peaks at time $4.610\sqrt{1024}\approx 147.5$ (as also seen in Fig. 7), the dashed red curve peaks at time $4.610\sqrt{2048}\approx 208.6$, and the dotted green curve peaks at time $4.610\sqrt{4096}\approx 295.0$.
@@ VI Conclusion
Previous research exploring search on the weighted truncated simplex lattice and linked complete graphs have shown that the weights between the cliques could be tuned to improve the runtime and success probability of quantum walk search algorithms by allowing probability to be freed from cliques. In these, the Laplacian and adjacency quantum walks searched in the same manner. In this paper, we explored the most restrictive means to join two cliques by investigating the weighted barbell graph of $N$ vertices, which has a single bridge of weight $w$. The Laplacian and adjacency quantum walks generally search this graph differently. For the Laplacian quantum walk, the weighted edge makes no difference; the evolution is the same regardless of the weight, where the probability is confined in each clique, and so the success probability reaches a maximum value of $1/2$ at time $\pi\sqrt{N}/2\sqrt{2}$. That is, the single bridge is too restrictive to affect the Laplacian quantum walk. For the adjacency quantum walk, most weights also make no difference, with the system evolving just like the Laplacian quantum walk. When $w=N/2$, however, the adjacency quantum walk does free some of the probability that was trapped in the unmarked clique, allowing the success probability to reach a value of 0.820 at time $2.518\sqrt{N}$, where the coefficient is the solution to a transcendental equation. Evolving beyond this, to a time of $3.265\sqrt{N}$, decreases the success probability, but causes 99.6% of the total probability to be in the marked clique. Then, the weight can be changed, for example to $w=1$, and the system will evolve to focus the success probability to the marked vertex, which takes an additional time of $1.345\sqrt{N}$. Altogether, this two-stage algorithm takes $4.610\sqrt{N}$ time to reach a success probability of $0.996$. Although this two-stage algorithm takes longer than search on the unweighted graph with repetitions, it could be useful in a scenario where it is expensive to repeat an algorithm, beyond the cost of querying the oracle.